One eighth of Emilia's coins are quarters and the rest are nickels. If the total value of her coins is $3.60, how many of each type of coin does she have?
Answer by fractalier(6550)

Call the numbers of quarters and nickels, q and n.
Since one eighth of Emilia's coins are quarters and the rest are nickels, we have n = 7q, or 7q - n = 0.
Our value equation is
25q + 5n = 360 (we work in cents)
Now multiply the top equation by five and add and get
25q + 5n = 360
+(35q - 5n = 0)
----------------
60q = 360
q = 6 quarters and then
n = 42 nickels